Please explain how to find the sum of two 16-bit integers. For example,

0 1 1 0 0 1 1 0 0 1 1 0 0 1 1 0  

0 1 0 1 0 1 0 1 0 1 0 1 0 1 0 1
Answer #1

Ex: 0+0=0

0+1=1

1+1=1 and carry 1

0 1 1 0 0 1 1 0 0 1 1 0 0 1 1 0 + 0 1 0 1 0 1 0 1 0 1 0 1 0 1 0 1= 0 1 0 1 1 1 0 1 1 1 0 1 1 1 0 1 1